OpenStudy (anonymous):

Write an expression for the volume of a cylinder with a height 7in. greater than the radius.

OpenStudy (anonymous):

V=pi*r*r*h is what i got but she toled me to right h in terms of r

OpenStudy (anonymous):

@bibby @Lakeekee_

OpenStudy (bibby):

write h in terms of r then

OpenStudy (anonymous):

but how

OpenStudy (bibby):

read the question

OpenStudy (bibby):

"a height 7in. greater than the radius. "

OpenStudy (anonymous):

so h=9

OpenStudy (anonymous):

@nerissa679

OpenStudy (anonymous):

@KamiBug

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

sorry bae i aint see yhu tagged me but iono this one sorry

OpenStudy (anonymous):

ok

OpenStudy (dan815):

let r be radius in V=pi*r*r*h and since h = r+7 V=pi*r*r*(r+7)

OpenStudy (dan815):

simplify

OpenStudy (anonymous):

omg dan i didnt know you was good at math

OpenStudy (dan815):